User name LeRose
Log entry time 11:50:03 on July28,2006
Entry number 175681
This entry is a followup to: 175598
keyword=Left Q2 OK
During this mornings controlled access techs increased the lead flow and turned down the lead heater on Left Q2. It now ramps up to 1600 and back down without any problem.